OpenStudy (anonymous):

The deposits Ginny makes at her bank each month form an arithmetic sequence. The deposit for month 3 is $150, and the deposit for month 5 is $180. Answer 1.what is common differences for each month? 2.write explicit formula for arithmetic sequence 3.amount in 12th month? 4.what month will deposit at least $500?

OpenStudy (anonymous):

\[T _{n}=a+(n-1)d\] \[T_{3}=a+2d\] \[150=a+2d\] \[T_{5}=a+4d\] \[180=a+4d[1]\] \[150=a+2d[2]\] \[[1]-[2]\] \[(180)-(150)=(a+4d)-(a+2d)\] \[30=2d\] \[d=?\]